### v1

# RankByGene: Gene-Guided Histopathology Representation Learning Through Cross-Modal Ranking Consistency

Abstract
Spatial transcriptomics (ST) provides essential spatial context by mapping gene expression within tissue, enabling detailed study of cellular heterogeneity and tissue organization. However, aligning ST data with histology images poses challenges due to inherent spatial distortions and modality-specific variations. Existing methods largely rely on direct alignment, which often fails to capture complex cross-modal relationships. To address these limitations, we propose a novel framework that aligns gene and image features using a ranking-based alignment loss, preserving relative similarity across modalities and enabling robust multi-scale alignment. To further enhance the alignment’s stability, we employ self-supervised knowledge distillation with a teacher-student network architecture, effectively mitigating disruptions from high dimensionality, sparsity, and noise in gene expression data. Extensive experiments on gene expression prediction and survival analysis demonstrate our framework’s effectiveness, showing improved alignment and predictive performance over existing methods and establishing a robust tool for gene-guided image representation learning in digital pathology.
1 Introduction
Digital pathology has advanced significantly in recent years, thanks to the availability of large number of digitized slides and rapid development of deep learning. Powerful learning methods have been proposed for image-based prediction at different scales, including individual cells [22, 5, 25], larger structures like glands [29, 69, 75], regions of interest [33, 20], and the whole slide [28, 34, 42, 32, 13]. However, these methods, despite their promise, are fundamentally limited by the amount of information one can extract from a digital slide. Most clinical outcomes are given at a whole slide-level. Due to the heterogeneity of the tissue microenvironment, it is very difficult for a machine learning algorithm to directly connect localized cell morphology and spatial context with clinical outcomes. Connecting cell morphology/context with their actual functionality is the missing link and can be the key to advancing image-based diagnosis and prognosis.
Transcriptomics [49, 72] reflects the genome-wide quantification of gene expression within a tissue, and these approaches provide a snapshot of gene activity within a variably heterogeneous collection of cells. Single-cell RNA sequencing [63, 62] attempts to address this cellular heterogeneity by profiling gene expression within individual level, enabling more precise cell typing and characterization. Despite its high resolution, this approach disrupts tissue architecture and loses the critical spatial context that defines an organized tissue [58].
Spatial transcriptomics (ST) bridges the gap by enabling gene expression profiling within tissue structures, preserving the spatial context of cellular activity [56]. It measures gene expression at specific locations, allowing for the mapping of cellular states and molecular interactions in their native micro-environments [61, 45]. This spatial dimension is crucial for studying the interactions between cells and their microenvironment, which are central to understanding disease dynamics and therapeutic responses [48, 19]. For example, ST can uncover molecular signatures unique to specific tumor regions, such as the core, invasive margins, and immune-infiltrated zones, enabling a nuanced view of tumor heterogeneity [31].
Figure 1: t-SNE visualization [66] of image representation from different methods. (a) Natural-Image-SSL. (b) Histopath-Image-SSL. (c) Gene-Image CL (d) Gene-Image CL with ranking loss. (e) RankByGene. We first use K-means clustering to assign a class label to each spot based on gene expression values. Then, we apply t-SNE to visualize the image features corresponding to each spot and color them according to their assigned class labels. A higher v-score [57] indicates that the image representation more precisely captures the distribution of gene expression values. Adding ranking loss (d) enhances the separability between image features of different classes. Our method (e), with both ranking loss and distillation shows the highest separability.
ST data provides new opportunities to enhance existing image-based prediction methods with novel biological information. Most ST data have both localized gene expressions and their corresponding histology image patches. One can exploit the two complementary modalities at a fine scale, unveiling the subtle visual cues from cell morphology and context that capture cell functionality. In this paper, we study the problem of learning gene-guided image representations through ST data. In particular, we use the gene expression to better align image features, so that these image features encode cell-functionality-specific visual information. As we will demonstrate in experiments, these gene-guided image features will improve image-based prediction of clinical outcomes. The learned features can be applied broadly to any image-only cohort, benefiting the digital pathology community.
Despite its scientific potential, multi-modal representation learning with ST data is challenging for various reasons. First and most importantly, it is fundamentally unknown how well the image and gene expressions could align. Different modalities may contain different information; ensuring complete alignment can be difficult or even impossible [27]. Secondly, the image and gene features are encoded with different networks and have different initial representation powers; while images can be encoded with CNNs and can be initialized with powerful foundational models [70, 14, 15, 74, 71, 68], typically gene feature encoding can only be based on fully-connected networks and trained from scratch [73, 30, 46]. Furthermore, the gene expression data is extremely high-dimensional (with more than 15,000 genes per sample) and very sparse (on average, 80% of gene expression values are zeros [30]). Finally, as a novel technology, available ST data often suffers from experimental artifacts, such as spatial dependent noise [6] and sparsity [35, 36].
To address these challenges, methods have been proposed to map both image patches and gene expressions into a shared latent feature space. Self-supervised contrastive learning losses have been adopted to ensure the image features and gene features from the same location (called a spot) are mapped together [30]. The standard contrastive loss, however, overlooks the potential similarity between different spots. Xie et al. [73] employ a smoothed variant of the CLIP loss [55] to encourage spots with similar image patches or similar gene expressions to be mapped closer in the feature space. These methods, however, do not fully address the aforementioned challenges and lack an effective global strategy to align multi-modal features.
In this paper, we propose a novel approach to align image and gene features across multiple scales, both locally and globally. The alignment method is designed to be robust to modality-specific distortions, addressing the challenges outlined earlier. Specifically, we introduce an innovative ranking-based alignment loss that captures the relative similarity ordering of feature representations across modalities. By enforcing consistency in cross-modal ranking, we ensure that the similarity relationships between gene expression features are reflected in the corresponding image features. This ranking-based approach offers several key advantages. First, it facilitates alignment across both local and global scales, enabling even distant features to interact and align. This global alignment complements the existing local matching of image and gene features of the same spot [73, 46]. Second, by focusing on matching rankings rather than exact similarity values, our method is more robust to distortions, allowing it to tolerate similarity variations especially between features that are far apart from one another.
To further enhance the robustness of the alignment process, we adopt a self-supervised knowledge distillation approach to stabilize the learning of gene-informed image representations. By utilizing a teacher-student network architecture with both weakly and strongly augmented image patches, we ensure that the student network effectively learns the aligned features. This approach helps the student network remain resilient to potential disruptions arising from imperfections in the gene expressions, such as high dimensionality, sparsity, and noisy or missing values.
Figure 1 shows the t-SNE visualization [66] of image features obtained from various methods: natural image-only self-supervised learning (SSL), histopathology image-only SSL, gene-image contrastive learning (CL), gene-image CL with ranking loss, and our method. Compared to other methods, the image features generated by our approach effectively capture the distribution of gene expression values.
In summary, we propose a novel multi-modal feature alignment method that integrates pathology images and ST data, enabling robust gene-informed image representation learning. Our contributions are three-fold:
  • We introduce a new cross-modal ranking consistency mechanism to enhance multi-modal alignment, achieving alignment at both local and global scales while remaining robust to distortions over long-range distances.
  • We apply a knowledge distillation approach to the multi-modal alignment task, further improving the robustness of the alignment process despite imperfections in gene expression features.
  • Extensive experiments across multiple downstream tasks, including gene expression prediction and survival prediction, demonstrate the superior performance of our framework compared to existing self-supervised alignment methods.
Figure 2: Overview of our RankbyGene framework. The framework begins with WSI Tiling, where WSIs are divided into tiles, each paired with a gene profile. In the feature extraction part, weak and strong augmentations of the image tiles are processed through a teacher encoder and student encoder, while a gene encoder extracts features from the gene profile. The feature alignment stage ensures that weakly and strongly augmented image features are aligned through intra-modal distillation loss and the image and gene features are aligned using gene-image contrastive loss. Meanwhile, our proposed cross-modal ranking consistency loss maintains consistent similarity ranking orders across two modalities.
2 Related Work
Histopathology Image Analysis. Histopathology image analysis is regarded as the gold standard for cancer diagnosis and treatment [41, 50, 9]. Due to the huge resolution of whole slide images (WSIs) and the difficulty of obtaining patch-level labels, WSI analysis is commonly performed in a weakly supervised setting [18], where only slide-level labels are available. Typical WSI analysis tasks include classification [28, 34, 77, 59, 54, 64, 78], which predicts the presence or absence of a tumor or identifies specific tumor subtypes, and survival prediction [13, 76, 81, 79, 60], which estimates patient mortality risk. Many multiple instance learning (MIL) [28, 34, 77, 53, 39, 82] based frameworks have been proposed to model relationships between patches more effectively, enabling more accurate slide-level predictions.
Image Feature Learning in Histopathology. The performance of WSI analysis heavily relies on image feature learning. With advancements in self-supervised learning (SSL) in computer vision, methods like SimCLR [16], MoCo [24], and DINO [11] have been adapted for image feature learning in histopathology, leading to the development of various foundation models pretrained on WSI, such as CTransPath [70], HIPT [14], and more recent models like UNI [15], GigaPath [74], CHIEF [71], and Virchow [68]. These foundation models have significantly improved WSI analysis performance. Additionally, task-specific fine-tuning [40, 37, 65] techniques have been introduced to further optimize image feature learning for downstream tasks.
Multi-Modal Pretraining in Histopathology. With the growing availability of multi-modal data and advancements in pretraining techniques across multiple modalities, CLIP-based [55, 8, 21, 38, 80] vision-language pretraining methods have been applied to text-image pathology datasets, resulting in the development of approaches like MI-zero [43], PLIP [26], and Conch [44].
Furthermore, the recent availability of ST data [30, 12] has expanded opportunities for multimodal learning in histopathology. More challenging tasks, such as gene expression prediction [30, 17, 12], have been introduced to evaluate the robustness of pretrained image encoders. In addition, several methods for image-gene pretraining, such as BLEEP [73] and mclSTExp [46], have also been proposed to enhance image feature learning. However, these approaches may not fully capture relationships between spots, nor adequately address the noise and sparsity in gene expression data. In this work, we focus on developing a robust multi-modal alignment approach that integrates image and gene features across multiple scales to achieve more effective gene-informed image representations.
3 Methods
Our learning pipeline is illustrated in Figure 2. The training data consists of pairs of image patches and gene expression profiles, with each pair corresponding to a specific tissue spot. Image patches are fed into image encoders, which include both a teacher and a student encoder to facilitate knowledge distillation for gene-guided image feature learning. This setup helps stabilize the image feature learning process during alignment with gene features. Meanwhile, gene expression data is processed by a gene encoder.
Both image and gene features are mapped into a shared latent feature space. To align these features, we employ a contrastive learning loss that ensures the image and gene features of the same tissue spot are closely matched. Additionally, we introduce a cross-modal ranking consistency loss to maintain the same relative ranking of similarity across spots, ensuring that for any given spot, its similarity to other spots is consistent in both image and gene feature spaces.
At inference time, we will only use the teacher image encoder to extract gene-guided image features from histopathology image patches. The feature can be used for different downstream tasks that make predictions based on histopathology images. We will first introduce the contrastive learning loss we use for spot-wise alignment, gene-image contrastive (Section 3.1). Next, we introduce the cross-modal ranking consistency loss (Section 3.2). We conclude this section with details on the intra-modal distillation for gene-guided image features (Section 3.3).
Preliminaries. Given a set of $N$ spots, $\{1,\ldots,N\}$, each with a pair of image patch and gene profile, we send them into two separate encoders (image encoder and gene encoder), mapping them into image features and gene features. $I=\{i_{1},i_{2},i_{3},\dots,i_{N}\}$ denotes the image features of the $N$ spots, and $G=\{g_{1},g_{2},g_{3},\dots,g_{N}\}$ their corresponding gene features.
3.1 Gene-Image Contrastive Loss
To align image and gene features, we first employ the InfoNCE Loss [51] , a commonly used loss function for aligning two modalities. InfoNCE encourages the model to pull positive pairs (image and gene features from the same spot) closer in the shared latent space while pushing apart negative pairs (image and gene features from different spots). The Gene-Image Contrastive loss is defined as follows:
$\mathcal{L}_{\text{gene-image}}=-\sum_{p=1}^{N}\left[\log\frac{\exp(\text{sim}(i_{p},g_{p})/\tau)}{\sum_{q=1}^{N}\exp(\text{sim}(i_{p},g_{q})/\tau)}\right]$ (1)
where $\text{sim}(i_{\ast},g_{\ast})$ is the cosine similarity between the image feature $i_{\ast}$ and the gene feature $g_{\ast}$. $\tau$ is a temperature parameter. This loss penalizes unmatched pairs by reducing their similarity while increasing the similarity between matched gene-image pairs.
3.2 Cross-Modal Ranking Consistency Loss
The InfoNCE loss ensures local alignment between image and gene features from the same tissue spot, but it does not address global alignment, which is essential for achieving more accurate and consistent cross-modal correspondences. Directly aligning distances between features from distant tissue spots is not practical, as long-range feature relationships may not be reliable. Instead, we propose that the relative ranking of distances between features is more robust and can provide a more trustworthy basis for alignment.
To leverage this idea, we introduce the Cross-Modal Ranking Consistency Loss. This loss function encourages the model to learn image representations while maintaining the relative similarity ordering of gene features across tissue spots. By focusing on the ranking of distances rather than exact alignments, the ranking loss facilitates a more reliable and robust global alignment. It complements the local alignment achieved by InfoNCE, while also capturing long-range interactions between features from different tissue spots. In doing so, the ranking consistency loss promotes a more comprehensive and stable cross-modal alignment, improving both local and global correspondences.
Figure 3: Illustration of the ranking loss intuition, when the gene features $g_{q}$ is closer to $g_{p}$ than $g_{r}$. Note the similarity is inverse proportional to the distance. Left: $i_{q}$ is also closer to $i_{p}$ than $i_{r}$, and furthermore, and the gap between image feature similarities $S_{p,q}^{I}$ and $S_{p,r}^{I}$ is bigger than the gap between gene feature similarities $S_{p,q}^{I}$ and $S_{p,r}^{I}$. This is the desirable case where $\ell(p,q,r)$ is negative. Middle: when the similarity ranking is the same between gene and image features, but when the gap between image feature similarities is smaller than the gap between gene feature similarities. $\ell(p,q,r)$ is positive, and will incur penalty. Right: when the similarity order is inconsistent, $\ell(p,q,r)$ is positive. Undesirable.
Our ranking loss is illustrated in Figure 3. For a spot of interest, $p$, we want to ensure the similarity between $p$ and two other spots, $q$ and $r$, are ranked consistently in both image feature and gene feature. For simplicity, we denote by $S^{I}_{p,q}=\text{sim}(i_{p},i_{q})$ and $S^{G}_{p,q}=\text{sim}(g_{p},g_{q})$ the similarity between $p$ and $q$’s image features and gene features. We also denote by $S^{I}_{p,r}$ and $S^{G}_{p,r}$ the image and gene feature similarity between $p$ and $r$. Assume $S^{G}_{p,q}>S^{G}_{p,r}$. In other words, $g_{q}$ is closer to $g_{p}$ compared with $g_{r}$. We would like to ensure the image features $i_{q}$ and $i_{r}$ also maintain the same order. Classic ordinal ranking loss [10] will just require the image feature similarity difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than a fixed positive value $\epsilon$. However, this may be undesirable especially for long interactions. To better preserve the relative positions between features, we require the difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than $S^{G}_{p,q}-S^{G}_{p,r}$. Formally, we define the function
$\ell(p,q,r)=\left(S^{G}_{p,q}-S^{G}_{p,r}\right)-\left(S^{I}_{p,q}-S^{I}_{p,r}\right)$ (2)
and intend to enforce $\ell(p,q,r)$ to be non-positive for all triplets $(p,q,r)$. Figure 3 illustrates different cases. In practice, we rewrite the function with the sign function
$\ell(p,q,r)=\text{sign}$ $(S^{G}_{p,q}-S^{G}_{p,r})\cdot$ (3)
$\left[\left(S^{G}_{p,q}-S^{G}_{p,r}\right)\right.\left.-\left(S^{I}_{p,q}-S^{I}_{p,r}\right)\right]$
This way, even if $S^{G}_{p,q}<S^{G}_{p,r}$, requiring $\ell(\cdot)$ to be non-positive will still enforce the relative ranking consistency between image and gene features of the three spots.
Finally, we need a loss to enforce $\ell(\cdot)$ to be nonnegative. To this end, we use a variant of the classic hinge loss, $\max\left\{0,\ell(p,q,r)\right\}$, which incurs penalty if and only if $\ell(p,q,r)$ is positive. Enumerating through all triplets, we have the cross-modal ranking consistency loss
$\mathcal{L}_{\text{rank}}=\sum_{p}\sum_{q\neq p}\sum_{r\neq p,q}\max\left\{0,\ell(p,q,r)\right\}$ (4)
Accelerating the Computation. During training, at each iteration, we will apply the loss to all spots within a mini-batch. $N$ becomes the batch size. However, this can still be computationally expensive, as it compares $O(N^{3})$ pairs of spots. To alleviate the burden, we still enumerate through all $p$’s. But for each $p$, instead of enumerating through all $O(N^{2})$ $(q,r)$ pairs, we only sample $O(N)$ random $(q,r)$ pairs. In particular, taking the list of spots except $p$, $L=(1,\ldots,p-1,p+1,\ldots,N)$. We randomly shuffle the sequence, getting a shuffled list, $L^{\prime}$. We go through the list, $L^{\prime}$, each time taking two consecutive spots as $q$ and $r$. For the last spot, we will pair it with the first spot in $L^{\prime}$. This gives us $N-1$ sample $(q,r)$ pairs, and ensures each spot appears in two of the sampled pairs. This reasonably covers similarity rankings of all the spots, but only using $O(N^{2})$ triplets.
3.3 Intra-Modal Distillation Loss
In our framework, we employ a teacher-student network architecture to achieve robust feature representations across differently augmented instances of the same pathology image, drawing on recent advances in self-supervised knowledge distillation for single-modality representation learning [11, 52]. To enhance stability and invariance in feature embeddings produced by the patch encoder, we apply both weak and strong augmentations to simulate the typical variability found in pathology images.
In this setup, the weakly augmented image is processed through the teacher encoder, while the strongly augmented version passes through the student encoder. The weights of the teacher encoder are incrementally updated using an Exponential Moving Average (EMA) of the student encoder’s weights, which helps stabilize the training. This strategy ensures that the student gradually learns stable features over time.
To enforce the consistency between the representations of the two augmented versions, we introduce the Image Consistency Loss:
$\mathcal{L}_{\text{distil}}=-\frac{1}{N}\sum_{p=1}^{N}\left[\log\frac{\exp(\text{sim}(i_{p}^{w},i_{p}^{s})/\tau)}{\sum_{q=1}^{N}\exp(\text{sim}(i_{p}^{w},i_{q}^{s})/\tau)}\right]$ (5)
where $N$ is the batch size, $i_{p}^{w}$ is the feature obtained from the weakly augmented image using the teacher image encoder, and $i_{p}^{s}$ is the feature from the strongly augmented image using the student image encoder. Minimizing this loss encourages the image encoder to learn representations resilient to such variations and potential disruptions from gene expression data characteristics, such as high dimensionality, sparsity, noise, and missing values.
3.4 Overall Loss Function
The total loss is a weighted sum of the Gene-Image Contrastive Loss, the Cross-Modal Ranking Loss, and the Intra-Modal Distillation Loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{gene-image}}+\lambda_{1}\mathcal{L}_{\text{rank}}+\lambda_{2}\mathcal{L}_{\text{distil}}$ (6)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that control the balance between the ranking consistency loss and the distillation loss, respectively.
4 Experiments and Results
4.1 Dataset
In this study, we use the breast and lung ST data from the HEST-1k dataset [30], which collects publicly available, high-quality ST datasets and applies standardized processing to all collected data. For WSIs, we utilize the TCGA-BRCA (Breast Invasive Carcinoma) and TCGA-LUAD (Lung Adenocarcinoma) datasets from TCGA cohorts [1]. Detailed descriptions of each dataset will be provided in the following sections.
Breast ST datasets. The Breast ST dataset used for training [7] contains 36 ST samples. Each sample contains approximately 15,000 genes, and after preprocessing, we obtained 13,620 spots from all ST samples. Following the external evaluation approach described in [17], we selected two Visium breast ST samples from 10x Genomics to evaluate our model’s performance in predicting gene expression. The first sample is the Human Breast Cancer: Visium Fresh Frozen, Whole Transcriptome [3], containing 4,898 spots. The second sample is the Human Breast Cancer (Block A Section 1) Visium [2], containing 3,813 spots. We refer to the first sample [3] as Breast-ST1 and the second sample [2] as Breast-ST2.
Lung ST datasets. The Lung ST dataset used for training originates from [47] and consists of 6 ST samples, each containing approximately 18, 000 genes, with a total of 12,891 spots. For testing, we select 4 lung ST samples from [67], which we refer to as Lung-ST, containing a total of 1,831 spots.
Cancer-Specific Gene List Selection To manage the high dimensionality of gene expression data, existing methods typically select a list of highly expressed genes for gene-image alignment, such as choosing the top 250 [17], top 100 [12], or top 50 [73, 30, 17] genes. However, this approach may overlook low-expression but significant genes, which can be crucial for downstream WSI analysis tasks. To help the model align images with cancer-specific genes, we select prognosis-related genes from the Human Protein Atlas [4]. Specifically, we first selected genes related to breast and lung cancer and then intersected these with genes in the ST data, resulting in final gene lists of 447 genes for breast ST and 1916 genes for lung ST data. We utilized these gene lists to train both our model and all multi-modal baseline methods. We will provide more details about the gene lists in the supplementary materials.
WSI Datasets. We use survival prediction on TCGA-BRCA (1, 022 WSIs) and TCGA-LUAD (515 WSIs) as our downstream task. Following the preprocessing steps in [42], we cut patches at $224\times 224$ pixels with 20×20$\times$20 × magnification.
4.2 Evaluation Metrics
For the gene expression prediction task, following previous methods [73, 17, 30], we use Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) as evaluation metrics to evaluate. For the survival prediction task, we adopt a 5-fold cross-validation setting as in [13, 76, 81] and use the cross-validated Concordance Index (C-Index) as the evaluation metric.
4.3 Implementation Details
A typical pipeline for our method consists of two stages. In the training stage, we first filter the genes using the selected cancer-related gene list, followed by multi-modal training with Breast ST [7] and Lung ST [47] dataset. In the testing stage, the trained image encoder extracts features for downstream tasks. The image encoder trained on Breast ST [7] dataset is evaluated on Breast-ST1, Breast-ST2, and TCGA-BRCA datasets, while the image encoder trained on Lung ST [47] dataset is evaluated on Lung-ST and TCGA-LUAD datasets.
Image Encoder Breast-ST1 Breast-ST2 Lung-ST
MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$ MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$ MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
ResNet-50 [23] 0.3633 0.2716 0.0867 0.4042 0.2982 0.1265 0.3105 0.2043 0.0345
CTransPath [70] 0.3572 0.2698 0.0983 0.3994 0.2957 0.1423 0.2945 0.1911 0.0395
UNI [15] 0.3559 0.2648 0.1063 0.4096 0.3178 0.1398 0.2731 0.1724 0.0432
HEST-FT [30] 0.3469 0.2552 0.1269 0.4071 0.3079 0.1592 0.2926 0.1874 0.0682
BLEEP [73] 0.3475 0.2571 0.1198 0.3982 0.2954 0.1437 0.2879 0.1806 0.0751
RankByGene 0.3412 0.2487 0.1803 0.3964 0.2927 0.1762 0.2662 0.1677 0.1019
Image Encoder C-Index
TCGA-BRCA TCGA-LUAD
ResNet-50 [23] 0.5859 $\pm$ 0.0521 0.5631 $\pm$ 0.0443
CTransPath [70] 0.6413 $\pm$ 0.0469 0.5742 $\pm$ 0.0363
UNI [15] 0.6684 $\pm$ 0.0432 0.5865 $\pm$ 0.0254
HEST-FT [30] 0.6531 $\pm$ 0.0453 0.5728 $\pm$ 0.0321
BLEEP [73] 0.6716 $\pm$ 0.0394 0.5778 $\pm$ 0.0383
RankByGene 0.6814 $\pm$ 0.0512 0.5945 $\pm$ 0.0293
Table 1: Results for the tasks of gene expression prediction and survival analysis using different patch encoders. We use Breast-ST1, Breast-ST2, and Lung-ST datasets for gene expression prediction. The slide-level prediction is performed using the ABMIL [28] framework.
We use UNI [15] as the backbone for the image encoder, followed by a 3-layer MLP as the projection head. A 3-layer MLP is also employed for the gene encoder, consistent with the previous method [73]. Both image and gene feature dimensions are set to 1024. For the image encoder, the learning rate is set to 0.0001, while for the gene encoder, it is set to 0.001. The other training settings and test setting are detailed in the supplementary materials.
4.4 Performance Comparison
We compare our method with image-only pretrained baselines, including ResNet-50 [23] (pretrained on ImageNet), CTransPath [70] and UNI [15] (both pretrained on pathology images), as well as gene-guided image feature learning baselines HEST-FT [30] and BLEEP [73].
Gene Prediction. As shown in Figure 4(a), RankByGene outperforms all baseline methods across the Breast-ST1, Breast-ST2, and Lung-ST datasets. Specifically, RankByGene achieves $11\%$ to $42\%$ improvement in PCC over the best baseline performance in each dataset. These results indicate that our method achieves stronger alignment between the image and gene expression data, with the obtained gene-informed image features effectively capturing the underlying global relationships between spots.
Survival Analysis. Using the models trained on Breast ST and Lung ST datasets, we evaluate their performance on survival prediction using TCGA-BRCA and TCGA-LUAD WSIs, respectively. Figure 4(b) presents the performance of different methods on these two datasets. We use the image encoder from each method to extract features on patches, then apply the ABMIL [28] framework to aggregate all patch-level features for slide-level prediction. The table shows that RankByGene achieves the highest C-Index of 0.6814 on the TCGA-BRCA dataset and 0.5945 on the TCGA-LUAD dataset compared to all baseline methods. The strong performance on these two datasets suggests that the improved image-gene alignment enables the model to capture survival-related signals more effectively, highlighting the potential of our method to enhance performance on challenging downstream tasks.
4.5 Ablation Study
$\lambda_{1}$ Breast-ST1
MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
0 0.3442 0.2516 0.1542
2 0.3431 0.2499 0.1776
5 0.3412 0.2487 0.1803
10 0.3427 0.2505 0.1762
$\lambda_{2}$ Breast-ST1
MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
0 0.3455 0.2535 0.1362
0.5 0.3438 0.2512 0.1509
1 0.3412 0.2487 0.1803
2 0.3436 0.2505 0.1544
$\mathcal{L}_{\text{rank}}$ $\mathcal{L}_{\text{distil}}$ Breast-ST1
MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
0.3469 0.2552 0.1269
0.3455 0.2535 0.1362
0.3442 0.2516 0.1542
0.3412 0.2487 0.1803
Encoder Breast-ST1
MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
ResNet50 0.3633 0.2716 0.0867
ResNet50 + Ours 0.3573 0.2682 0.1035
UNI 0.3559 0.2648 0.1063
UNI + Ours 0.3412 0.2487 0.1803
Table 2: RankByGene ablation experiments. We conduct all ablation experiments on the gene expression prediction task and the Breast-ST1 dataset.
Weight of Cross-Modal Ranking Consistency Loss. We compare the effect of different values for the weight of the ranking consistency loss $\lambda_{1}$ on the Breast-ST dataset. The results in Figure 5(a) show that the model is robust to the choice of the hyperparameter. The performance across all the tested values $\lambda_{1}>0$ significantly outperforms the baselines, with the best result using $\lambda_{1}=5$.
Weight of Intra-Modal Distillation Loss. Figure 5(b) presents the ablation study results for the image consistency loss weight $\lambda_{2}$ on the Breast-ST1 dataset. Our method shows stable performance across MAE, MSE, and PCC at various values of $\lambda_{2}$, with consistent improvements over the baseline methods. These results indicate the benefits of incorporating image consistency loss, as it further enhances the robustness of image feature learning during alignment with gene features.
Ablation Study on Loss Component. Figure 5(c) presents the ablation study results on the impact of individual loss components on the Breast-ST1 dataset. When both loss components are removed, the model exhibits the lowest PCC and higher MAE and MSE values. Including either $\mathcal{L}_{\text{rank}}$ or $\mathcal{L}_{\text{distil}}$ enhances performance, with the best results achieved when both components are applied. The results indicate that both the ranking loss and the image consistency loss contribute significantly to the model’s performance.
Figure 5(d) presents the results of applying our method to different image encoders on the Breast ST1 dataset, specifically ResNet50 (a CNN-based model pretrained on ImageNet) and UNI (a ViT-based model pretrained on pathology images using the DINO framework [11]). The results demonstrate that our method consistently enhances the performance of each encoder in the gene prediction task, as evidenced by lower MAE and MSE values and higher PCC scores. This finding shows the adaptability of our approach, demonstrating its effectiveness across various encoder architectures.
4.6 Qualitative Result
Figure 5: Visualization of PTGES3 gene expression predictions from different methods, with all values normalized to the range of 0 to 1. Compared to the baseline, our predictions show the closest match to the ground truth.
Rank Accuracy During Training. We define Rank Accuracy as a metric to evaluate whether the ranking relationships among gene features are effectively captured by the image features during training. After the end of each epoch, Rank Accuracy is calculated as follows: the first gene-image pair in each batch is randomly selected as the anchor pair, along with two additional target pairs. Rank Accuracy is determined by checking whether the target pair with higher gene feature similarity to the anchor pair also has a higher image feature similarity. This process is repeated eight times to calculate the final Rank Accuracy. In Figure 6, we compare the change of the rank accuracy during training between different methods. The figure shows that our method achieves a larger improvement in accuracy during training compared to the baseline methods, suggesting that the rank relationships in the gene features are effectively captured by image features during the training process.
Figure 6: Rank accuracy for different methods during training.
Visualization of Cancer Marker Gene. To evaluate the accuracy of our model’s predictions for critical genes, we selected PTGES3 as the marker gene for visualization on the Breast-ST1 dataset due to its association with poor prognosis in breast invasive carcinoma across two independent cohorts [4]. We visualized the gene prediction values for the PTGES3 from different methods across each spot, including ground truth values and the corresponding WSI region, as shown in Figure 5. Compared to the baseline method, our approach provides more accurate predictions for the PTGES3, demonstrating its effectiveness in capturing important gene expression patterns. We also provide visualizations of other marker genes for lung cancer in the supplementary materials.
5 Conclusion
Our proposed framework establishes a robust and scalable approach for efficiently aligning spatial transcriptomics data with histopathology images, overcoming key challenges in cross-modal alignment. We propose a ranking-based mechanism to capture similarity relationships between gene and image features, facilitating alignment that is resilient to distortions and effective across both local and global scales. This approach preserves essential spatial and molecular patterns, which is critical for accurately capturing the biological complexity of tissue structures. Additionally, our use of a self-supervised knowledge distillation method further enhances the alignment process, enabling the model to adaptively handle the high dimensionality, sparsity, and noise inherent in gene expression data. Through comprehensive experiments, we demonstrate our framework not only improves alignment quality but also enhances predictive capabilities in gene expression prediction and survival analysis. Ultimately, this work provides a foundational tool for multi-modal representation learning in digital pathology.

### v2

@@
 To further enhance the alignment’s stability, we employ self-supervised knowledge distillation with a teacher-student network architecture, effectively mitigating disruptions from high dimensionality, sparsity, and noise in gene expression data.
-Extensive experiments on gene expression prediction and survival analysis demonstrate our framework’s effectiveness, showing improved alignment and predictive performance over existing methods and establishing a robust tool for gene-guided image representation learning in digital pathology.
+Extensive experiments on seven public datasets that encompass gene expression prediction, slide-level classification, and survival analysis demonstrate the efficacy of our method, showing improved alignment and predictive performance over existing methods.
 1 Introduction
@@
 Powerful learning methods have been proposed for image-based prediction at different scales, including individual cells [23, 5, 27], larger structures like glands [31, 73, 80], regions of interest [35, 21], and the whole slide [30, 36, 44, 34, 13].
-However, these methods, despite their promise, are fundamentally limited by the amount of information one can extract from a digital slide.
-Most clinical outcomes are given at a whole slide-level.
-Due to the heterogeneity of the tissue microenvironment, it is very difficult for a machine learning algorithm to directly connect localized cell morphology and spatial context with clinical outcomes.
-Connecting cell morphology/context with their actual functionality is the missing link and can be the key to advancing image-based diagnosis and prognosis.
-Transcriptomics [49, 72] reflects the genome-wide quantification of gene expression within a tissue, and these approaches provide a snapshot of gene activity within a variably heterogeneous collection of cells.
-Single-cell RNA sequencing [63, 62] attempts to address this cellular heterogeneity by profiling gene expression within individual level, enabling more precise cell typing and characterization.
-Despite its high resolution, this approach disrupts tissue architecture and loses the critical spatial context that defines an organized tissue [58].
-Spatial transcriptomics (ST) bridges the gap by enabling gene expression profiling within tissue structures, preserving the spatial context of cellular activity [56].
-It measures gene expression at specific locations, allowing for the mapping of cellular states and molecular interactions in their native micro-environments [61, 45].
-This spatial dimension is crucial for studying the interactions between cells and their microenvironment, which are central to understanding disease dynamics and therapeutic responses [48, 19].
-For example, ST can uncover molecular signatures unique to specific tumor regions, such as the core, invasive margins, and immune-infiltrated zones, enabling a nuanced view of tumor heterogeneity [31].
-Figure 1: t-SNE visualization [66] of image representation from different methods.
-(a) Natural-Image-SSL.
-(b) Histopath-Image-SSL.
-(c) Gene-Image CL (d) Gene-Image CL with ranking loss.
-(e) RankByGene.
-We first use K-means clustering to assign a class label to each spot based on gene expression values.
-Then, we apply t-SNE to visualize the image features corresponding to each spot and color them according to their assigned class labels.
-A higher v-score [57] indicates that the image representation more precisely captures the distribution of gene expression values.
-Adding ranking loss (d) enhances the separability between image features of different classes.
-Our method (e), with both ranking loss and distillation shows the highest separability.
-ST data provides new opportunities to enhance existing image-based prediction methods with novel biological information.
+However, despite their promise, these methods are inherently constrained by the information that can be extracted from a digital slide.
+Bridging the gap between cell morphology, contextual information, and actual functionality remains a critical challenge, one that could be key to advancing image-based diagnosis and prognosis.
+Figure 1: t-SNE [70] visualization of image features of different spots in a ST slide.
+We show the feature learned using different methods, including (a) SSL on natural images, (b) SSL on histopathology images, (c) CL on ST data using InfoNCE loss [32], and (d) RankByGene.
+Learning with gene information (c) clearly outperforms learning with images alone ((a) and (b)).
+Furthermore, our method (d) achieves even greater improvement, demonstrating better inter-class separability than (c), thanks to the proposed contributions.
+Each spot is assigned a label via K-means clustering on gene expression values; the corresponding image features are then color-coded by these labels.
+We also provide quantitative measures using v-score [61], where a higher value indicates better alignment of image features with gene expression.
+Transcriptomics [51, 76] quantifies gene expression across tissues, providing a snapshot of cellular heterogeneity.
+Single-cell RNA sequencing [67, 66] refines this by profiling individual cells, enabling precise cell characterization but disrupting tissue architecture and losing spatial context [62].
+Spatial transcriptomics (ST) overcomes this limitation by preserving spatial organization while mapping gene expression [60].
+It captures cellular states and molecular interactions within their native environment [65, 47], crucial for understanding disease dynamics and therapeutic responses [50, 20].
+For instance, ST reveals distinct molecular signatures in tumor regions, such as the core, invasive margins, and immune-infiltrated zones, offering deeper insights into tumor heterogeneity [33].
+However, its high cost remains a significant barrier to widespread clinical adoption.
+Despite the high cost, ST data provides new opportunities to enhance existing affordable image-based prediction methods with novel biological information.
 Most ST data have both localized gene expressions and their corresponding histology image patches.
 One can exploit the two complementary modalities at a fine scale, unveiling the subtle visual cues from cell morphology and context that capture cell functionality.
+As shown in recent studies [12, 32, 48, 77], visual information including the density and spatial organization of various cell types (e.g., epithelial, lymphocyte, fibroblast) correlates with gene pathways and expressions.
+If correctly distilled, these gene-correlated visual cues can be valuable for downstream tasks.
 In this paper, we study the problem of learning gene-guided image representations through ST data.
@@
 [77] employ a smoothed variant of the CLIP loss [59] to encourage spots with similar image patches or similar gene expressions to be mapped closer in the feature space.
+Similarly, Min et al.
+[48] adopt a similar approach to enhance feature alignment.
 These methods, however, do not fully address the aforementioned challenges and lack an effective global strategy to align multi-modal features.
-In this paper, we propose a novel approach to align image and gene features across multiple scales, both locally and globally.
+RankByGene.
+We propose RankByGene, a novel approach to align image and gene features across multiple scales, both locally and globally.
 The alignment method is designed to be robust to modality-specific distortions, addressing the challenges outlined earlier.
-Specifically, we introduce an innovative ranking-based alignment loss that captures the relative similarity ordering of feature representations across modalities.
-By enforcing consistency in cross-modal ranking, we ensure that the similarity relationships between gene expression features are reflected in the corresponding image features.
+Specifically, we introduce an innovative ranking-based alignment loss ensuring that similarity relationships between gene expression features are reflected in the corresponding image features.
 This ranking-based approach offers several key advantages.
@@
 This global alignment complements the existing local matching of image and gene features of the same spot [77, 48].
-Second, by focusing on matching rankings rather than exact similarity values, our method is more robust to distortions, allowing it to tolerate similarity variations especially between features that are far apart from one another.
-To further enhance the robustness of the alignment process, we adopt a self-supervised knowledge distillation approach to stabilize the learning of gene-informed image representations.
+Second, by focusing on matching rankings rather than exact similarity values, our method is more robust to distortions, allowing it to tolerate similarity variations, especially between features that are far apart from one another.
+To further enhance the robustness of the alignment, we adopt a self-supervised knowledge distillation approach to stabilize the learning of gene-informed image representations.
 By utilizing a teacher-student network architecture with both weakly and strongly augmented image patches, we ensure that the student network effectively learns the aligned features.
-This approach helps the student network remain resilient to potential disruptions arising from imperfections in the gene expressions, such as high dimensionality, sparsity, and noisy or missing values.
-Figure 1 shows the t-SNE visualization [66] of image features obtained from various methods: natural image-only self-supervised learning (SSL), histopathology image-only SSL, gene-image contrastive learning (CL), gene-image CL with ranking loss, and our method.
-Compared to other methods, the image features generated by our approach effectively capture the distribution of gene expression values.
+This approach helps the student network remain resilient to potential disruptions arising from imperfections in ST gene expression measurements, such as high dimensionality, sparsity, and noisy or missing values.
+The strength of our method is demonstrated in Figure 1, which shows the t-SNE visualization [70] of image features from various methods: natural image-only self-supervised learning (SSL), histopathology image-only SSL, gene-image contrastive learning (CL) using InfoNCE [32], and RankByGene.
+Gene-image CL clearly obtains image features better aligned with gene expressions.
+Our approach further boosts the features using the proposed contributions.
 In summary, we propose a novel multi-modal feature alignment method that integrates pathology images and ST data, enabling robust gene-informed image representation learning.
-Our contributions are three-fold:
-  • We introduce a new cross-modal ranking consistency mechanism to enhance multi-modal alignment, achieving alignment at both local and global scales while remaining robust to distortions over long-range distances.
-  • We apply a knowledge distillation approach to the multi-modal alignment task, further improving the robustness of the alignment process despite imperfections in gene expression features.
-  • Extensive experiments across multiple downstream tasks, including gene expression prediction and survival prediction, demonstrate the superior performance of our framework compared to existing self-supervised alignment methods.
+Our contributions are threefold:
+  • We propose a novel cross-modal ranking consistency mechanism that enhances gene-image alignment at both local and global scales, while maintaining robustness to long-range distortions.
+  • We introduce a knowledge distillation module to the gene-image alignment framework, further improving the robustness of the alignment against artifacts in ST data gene expression such as sparsity and noisy measurements.
+  • Extensive experiments on gene expression prediction, gene-related slide-level classification and survival analysis tasks show that our method outperforms existing ones, underscoring its robustness and efficacy.
 Figure 2: Overview of our RankbyGene framework.
-The framework begins with WSI Tiling, where WSIs are divided into tiles, each paired with a gene profile.
-In the feature extraction part, weak and strong augmentations of the image tiles are processed through a teacher encoder and student encoder, while a gene encoder extracts features from the gene profile.
+The framework begins with WSI Tiling, where WSIs are cut into patches, each paired with a gene spot.
+In feature extraction, weak and strong augmentations of the patches are processed through a teacher and student encoder, while a gene encoder extracts features from the gene profile.
 The feature alignment stage ensures that weakly and strongly augmented image features are aligned through intra-modal distillation loss and the image and gene features are aligned using gene-image contrastive loss.
-Meanwhile, our proposed cross-modal ranking consistency loss maintains consistent similarity ranking orders across two modalities.
+Meanwhile, our proposed cross-modal ranking consistency loss maintains consistent similarity ranking across two modalities.
 2 Related Work
@@
 Typical WSI analysis tasks include classification [30, 36, 82, 63, 58, 68, 83], which predicts the presence or absence of a tumor or identifies specific tumor subtypes, and survival prediction [13, 81, 86, 84, 64], which estimates patient mortality risk.
-Many multiple instance learning (MIL) [28, 34, 77, 53, 39, 82] based frameworks have been proposed to model relationships between patches more effectively, enabling more accurate slide-level predictions.
+Multiple instance learning (MIL) [30, 36, 82, 57, 41, 87] based frameworks have been proposed to model relationships between patches more effectively, enabling more accurate slide-level predictions.
 Image Feature Learning in Histopathology.
@@
 The training data consists of pairs of image patches and gene expression profiles, with each pair corresponding to a specific tissue spot.
-Image patches are fed into image encoders, which include both a teacher and a student encoder to facilitate knowledge distillation for gene-guided image feature learning.
+Image patches are fed into image encoders that include both a teacher and a student encoder to facilitate knowledge distillation for gene-guided image feature learning.
 This setup helps stabilize the image feature learning process during alignment with gene features.
-Meanwhile, gene expression data is processed by a gene encoder.
+Meanwhile, gene expression data are processed by a gene encoder.
 Both image and gene features are mapped into a shared latent feature space.
 To align these features, we employ a contrastive learning loss that ensures the image and gene features of the same tissue spot are closely matched.
-Additionally, we introduce a cross-modal ranking consistency loss to maintain the same relative ranking of similarity across spots, ensuring that for any given spot, its similarity to other spots is consistent in both image and gene feature spaces.
+A main contribution is that we introduce a cross-modal ranking consistency loss, ensuring that for any given spot, its similarity to other spots is consistent in both image and gene feature spaces.
+This ranking loss ensures robust global alignment even for spots whose features are not close to each other.
 At inference time, we will only use the teacher image encoder to extract gene-guided image features from histopathology image patches.
@@
 Undesirable.
-Our ranking loss is illustrated in Figure 3.
-For a spot of interest, $p$, we want to ensure the similarity between $p$ and two other spots, $q$ and $r$, are ranked consistently in both image feature and gene feature.
+Our ranking loss is inspired by the classic ordinal ranking loss [10], but is modified to better address challenges in our gene-image alignment task.
+As illustrated in Figure 3, for a spot of interest, $p$, we want to ensure the similarity between $p$ and two other spots, $q$ and $r$, are ranked consistently in both image feature and gene feature.
 For simplicity, we denote by $S^{I}_{p,q}=\text{sim}(i_{p},i_{q})$ and $S^{G}_{p,q}=\text{sim}(g_{p},g_{q})$ the similarity between $p$ and $q$’s image features and gene features.
@@
 Classic ordinal ranking loss [10] will just require the image feature similarity difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than a fixed positive value $\epsilon$.
-However, this may be undesirable especially for long interactions.
-To better preserve the relative positions between features, we require the difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than $S^{G}_{p,q}-S^{G}_{p,r}$.
-Formally, we define the function
+In practice, however, we often notice that this loss does not work effectively; image features are often more tightly packed compared with gene features, thus just maintaining the ranking of image similarities will not making much changes, at least in the early alignment stages.
+To further accelerate the alignment, we propose a modified ranking loss that pushes image features further apart if possible.
+In our new ranking loss, we forcing similarity among images to be bigger than similarity among corresponding genes; it pushes image features apart while preserving the ranking.
+Formally, we require the difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than $S^{G}_{p,q}-S^{G}_{p,r}$.
+We define
 $\ell(p,q,r)=\left(S^{G}_{p,q}-S^{G}_{p,r}\right)-\left(S^{I}_{p,q}-S^{I}_{p,r}\right)$ (2)
@@
 Accelerating the Computation.
-During training, at each iteration, we will apply the loss to all spots within a mini-batch.
-$N$ becomes the batch size.
+During training, at each iteration, we will apply the loss to all spots within a mini-batch (batch size $N$).
 However, this can still be computationally expensive, as it compares $O(N^{3})$ pairs of spots.
@@
 This reasonably covers similarity rankings of all the spots, but only using $O(N^{2})$ triplets.
+Image Encoder Breast-ST1 Breast-ST2 Lung-ST
+MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$ MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$ MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
+(a) Results for gene expression prediction (top 250 genes among all genes)
+ResNet-50 [25] 0.494 $\pm$ 0.004 0.401 $\pm$ 0.006 0.069 $\pm$ 0.004 0.465 $\pm$ 0.003 0.384 $\pm$ 0.008 0.069 $\pm$ 0.008 0.709 $\pm$ 0.013 0.695 $\pm$ 0.019 0.026 $\pm$ 0.009
+CTransPath [74] 0.489 $\pm$ 0.005 0.388 $\pm$ 0.012 0.088 $\pm$ 0.007 0.459 $\pm$ 0.012 0.352 $\pm$ 0.013 0.132 $\pm$ 0.009 0.690 $\pm$ 0.012 0.635 $\pm$ 0.017 0.039 $\pm$ 0.005
+UNI [15] 0.488 $\pm$ 0.008 0.383 $\pm$ 0.010 0.096 $\pm$ 0.008 0.461 $\pm$ 0.015 0.335 $\pm$ 0.017 0.154 $\pm$ 0.014 0.681 $\pm$ 0.011 0.619 $\pm$ 0.019 0.043 $\pm$ 0.007
+ST-Net [24] 0.491 $\pm$ 0.004 0.395 $\pm$ 0.013 0.085 $\pm$ 0.009 0.464 $\pm$ 0.012 0.362 $\pm$ 0.009 0.104 $\pm$ 0.011 0.685 $\pm$ 0.015 0.615 $\pm$ 0.014 0.048 $\pm$ 0.006
+HisToGene [56] 0.485 $\pm$ 0.006 0.385 $\pm$ 0.017 0.102 $\pm$ 0.011 0.458 $\pm$ 0.015 0.329 $\pm$ 0.015 0.144 $\pm$ 0.007 0.674 $\pm$ 0.016 0.603 $\pm$ 0.015 0.055 $\pm$ 0.004
+HEST-FT [32] 0.482 $\pm$ 0.011 0.379 $\pm$ 0.015 0.125 $\pm$ 0.013 0.452 $\pm$ 0.013 0.324 $\pm$ 0.016 0.207 $\pm$ 0.005 0.661 $\pm$ 0.011 0.594 $\pm$ 0.018 0.078 $\pm$ 0.004
+BLEEP [77] 0.477 $\pm$ 0.015 0.378 $\pm$ 0.017 0.138 $\pm$ 0.009 0.454 $\pm$ 0.016 0.326 $\pm$ 0.011 0.212 $\pm$ 0.006 0.669 $\pm$ 0.015 0.589 $\pm$ 0.012 0.081 $\pm$ 0.001
+RankByGene 0.472 $\pm$ 0.010 0.371 $\pm$ 0.012 0.185 $\pm$ 0.005 0.447 $\pm$ 0.008 0.318 $\pm$ 0.010 0.231 $\pm$ 0.012 0.648 $\pm$ 0.013 0.570 $\pm$ 0.011 0.097 $\pm$ 0.006
+(b) Results for gene expression prediction (top 250 genes in cancer-specific gene list)
+ResNet-50 [25] 0.381 $\pm$ 0.005 0.291 $\pm$ 0.008 0.089 $\pm$ 0.008 0.432 $\pm$ 0.006 0.336 $\pm$ 0.005 0.112 $\pm$ 0.011 0.317 $\pm$ 0.006 0.215 $\pm$ 0.007 0.041 $\pm$ 0.012
+CTransPath [74] 0.375 $\pm$ 0.007 0.287 $\pm$ 0.005 0.098 $\pm$ 0.012 0.429 $\pm$ 0.004 0.329 $\pm$ 0.007 0.132 $\pm$ 0.015 0.312 $\pm$ 0.006 0.203 $\pm$ 0.009 0.047 $\pm$ 0.014
+UNI [15] 0.369 $\pm$ 0.004 0.282 $\pm$ 0.009 0.116 $\pm$ 0.014 0.438 $\pm$ 0.007 0.331 $\pm$ 0.005 0.127 $\pm$ 0.014 0.288 $\pm$ 0.004 0.183 $\pm$ 0.005 0.054 $\pm$ 0.009
+ST-Net [24] 0.372 $\pm$ 0.007 0.284 $\pm$ 0.006 0.108 $\pm$ 0.011 0.426 $\pm$ 0.006 0.327 $\pm$ 0.012 0.125 $\pm$ 0.012 0.309 $\pm$ 0.007 0.198 $\pm$ 0.011 0.057 $\pm$ 0.010
+HisToGene [56] 0.366 $\pm$ 0.009 0.279 $\pm$ 0.011 0.134 $\pm$ 0.014 0.422 $\pm$ 0.005 0.318 $\pm$ 0.016 0.146 $\pm$ 0.015 0.306 $\pm$ 0.012 0.194 $\pm$ 0.010 0.064 $\pm$ 0.011
+HEST-FT [32] 0.362 $\pm$ 0.010 0.277 $\pm$ 0.011 0.142 $\pm$ 0.011 0.417 $\pm$ 0.006 0.313 $\pm$ 0.013 0.157 $\pm$ 0.019 0.301 $\pm$ 0.008 0.188 $\pm$ 0.007 0.077 $\pm$ 0.007
+BLEEP [77] 0.363 $\pm$ 0.008 0.276 $\pm$ 0.014 0.138 $\pm$ 0.013 0.415 $\pm$ 0.007 0.311 $\pm$ 0.015 0.160 $\pm$ 0.009 0.297 $\pm$ 0.009 0.187 $\pm$ 0.009 0.084 $\pm$ 0.005
+RankByGene 0.354 $\pm$ 0.007 0.272 $\pm$ 0.009 0.187 $\pm$ 0.010 0.407 $\pm$ 0.006 0.306 $\pm$ 0.012 0.172 $\pm$ 0.008 0.276 $\pm$ 0.008 0.176 $\pm$ 0.005 0.103 $\pm$ 0.004
+Table 1: Gene expression prediction results: (a) top 250 genes and (b) top 250 cancer-specific genes.
 3.3 Intra-Modal Distillation Loss
@@
 4 Experiments and Results
-4.1 Dataset
-In this study, we use the breast and lung ST data from the HEST-1k dataset [30], which collects publicly available, high-quality ST datasets and applies standardized processing to all collected data.
-For WSIs, we utilize the TCGA-BRCA (Breast Invasive Carcinoma) and TCGA-LUAD (Lung Adenocarcinoma) datasets from TCGA cohorts [1].
-Detailed descriptions of each dataset will be provided in the following sections.
-Breast ST datasets.
-The Breast ST dataset used for training [7] contains 36 ST samples.
-Each sample contains approximately 15,000 genes, and after preprocessing, we obtained 13,620 spots from all ST samples.
-Following the external evaluation approach described in [17], we selected two Visium breast ST samples from 10x Genomics to evaluate our model’s performance in predicting gene expression.
-The first sample is the Human Breast Cancer: Visium Fresh Frozen, Whole Transcriptome [3], containing 4,898 spots.
-The second sample is the Human Breast Cancer (Block A Section 1) Visium [2], containing 3,813 spots.
-We refer to the first sample [3] as Breast-ST1 and the second sample [2] as Breast-ST2.
-Lung ST datasets.
-The Lung ST dataset used for training originates from [47] and consists of 6 ST samples, each containing approximately 18, 000 genes, with a total of 12,891 spots.
-For testing, we select 4 lung ST samples from [67], which we refer to as Lung-ST, containing a total of 1,831 spots.
-Cancer-Specific Gene List Selection To manage the high dimensionality of gene expression data, existing methods typically select a list of highly expressed genes for gene-image alignment, such as choosing the top 250 [17], top 100 [12], or top 50 [73, 30, 17] genes.
-However, this approach may overlook low-expression but significant genes, which can be crucial for downstream WSI analysis tasks.
-To help the model align images with cancer-specific genes, we select prognosis-related genes from the Human Protein Atlas [4].
-Specifically, we first selected genes related to breast and lung cancer and then intersected these with genes in the ST data, resulting in final gene lists of 447 genes for breast ST and 1916 genes for lung ST data.
-We utilized these gene lists to train both our model and all multi-modal baseline methods.
-We will provide more details about the gene lists in the supplementary materials.
-WSI Datasets.
-We use survival prediction on TCGA-BRCA (1, 022 WSIs) and TCGA-LUAD (515 WSIs) as our downstream task.
-Following the preprocessing steps in [42], we cut patches at $224\times 224$ pixels with 20×20$\times$20 × magnification.
-4.2 Evaluation Metrics
-For the gene expression prediction task, following previous methods [73, 17, 30], we use Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) as evaluation metrics to evaluate.
-For the survival prediction task, we adopt a 5-fold cross-validation setting as in [13, 76, 81] and use the cross-validated Concordance Index (C-Index) as the evaluation metric.
-4.3 Implementation Details
-A typical pipeline for our method consists of two stages.
-In the training stage, we first filter the genes using the selected cancer-related gene list, followed by multi-modal training with Breast ST [7] and Lung ST [47] dataset.
-In the testing stage, the trained image encoder extracts features for downstream tasks.
-The image encoder trained on Breast ST [7] dataset is evaluated on Breast-ST1, Breast-ST2, and TCGA-BRCA datasets, while the image encoder trained on Lung ST [47] dataset is evaluated on Lung-ST and TCGA-LUAD datasets.
-Image Encoder Breast-ST1 Breast-ST2 Lung-ST
-MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$ MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$ MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
-ResNet-50 [23] 0.3633 0.2716 0.0867 0.4042 0.2982 0.1265 0.3105 0.2043 0.0345
-CTransPath [70] 0.3572 0.2698 0.0983 0.3994 0.2957 0.1423 0.2945 0.1911 0.0395
-UNI [15] 0.3559 0.2648 0.1063 0.4096 0.3178 0.1398 0.2731 0.1724 0.0432
-HEST-FT [30] 0.3469 0.2552 0.1269 0.4071 0.3079 0.1592 0.2926 0.1874 0.0682
-BLEEP [73] 0.3475 0.2571 0.1198 0.3982 0.2954 0.1437 0.2879 0.1806 0.0751
-RankByGene 0.3412 0.2487 0.1803 0.3964 0.2927 0.1762 0.2662 0.1677 0.1019
+We evaluate our method on gene expression prediction, and downstream tasks including gene mutation prediction, receptor status classification, and survival analysis.
+Datasets and preprocessing.
+We use the breast and lung ST data from the HEST-1k dataset [32], which collects publicly available, high-quality ST datasets and applies standardized processing to all collected data.
+For WSIs, we utilize datasets from TCGA cohorts [1] and BCNB [78] for classification and survival prediction tasks.
+Further details of the datasets are provided in the supplementary.
+Following [32], we apply a three-step preprocessing pipeline: L1 normalization, log transformation, and 8-neighborhood smoothing to all ST data.
+The processed data is used in both training and testing.
+Baselines.
+We compare our method with image-only pretrained baselines, including ResNet-50 [25] (pretrained on ImageNet), CTransPath [74], and UNI [15] (all pretrained on pathology images), as well as gene-guided image feature learning baselines HEST-FT [32] and BLEEP [77].
+Since HEST-FT and BLEEP are both gene-image pretraining methods, for a fair comparison we used UNI for the image encoder initialization and use the original initialization for the other baselines.
+We also compare our method with methods that are specifically designed for gene expression prediction, such as ST-Net [24] and HisToGene [56], using their default initialization.
+Evaluation metrics.
+For the gene expression prediction task, we adopt Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) as evaluation metrics, following previous studies [77, 17, 32].
+The WSI classification performance is assessed using the Area Under the Receiver Operating Characteristic Curve (AUC).
+For survival prediction, we follow [13, 81, 86] and evaluate the model using the Concordance Index (C-Index).
+Image Encoder BCNB TCGA-LUAD mutation
+BR PR HER2 EGFR KRAS STK11 TP53
+ResNet-50 [25] 0.833 ± 0.039 0.785 ± 0.032 0.709 ± 0.031 0.729 ± 0.068 0.643 ± 0.048 0.739 ± 0.015 0.691 ± 0.023
+CTransPath [74] 0.857 ± 0.023 0.792 ± 0.015 0.716 ± 0.019 0.812 ± 0.024 0.682 ± 0.059 0.797 ± 0.025 0.783 ± 0.022
+UNI [15] 0.891 ± 0.027 0.809 ± 0.018 0.724 ± 0.032 0.845 ± 0.047 0.712 ± 0.046 0.843 ± 0.021 0.827 ± 0.019
+HEST-FT [32] 0.901 ± 0.021 0.811 ± 0.013 0.721 ± 0.015 0.851 ± 0.036 0.716 ± 0.051 0.851 ± 0.028 0.834 ± 0.026
+BLEEP [77] 0.899 ± 0.042 0.817 ± 0.019 0.734 ± 0.024 0.862 ± 0.028 0.713 ± 0.049 0.855 ± 0.019 0.838 ± 0.029
+RankByGene 0.915 ± 0.024 0.829 ± 0.012 0.759 ± 0.028 0.855 ± 0.037 0.729 ± 0.065 0.861 ± 0.024 0.845 ± 0.033
 Image Encoder C-Index
 TCGA-BRCA TCGA-LUAD
-ResNet-50 [23] 0.5859 $\pm$ 0.0521 0.5631 $\pm$ 0.0443
-CTransPath [70] 0.6413 $\pm$ 0.0469 0.5742 $\pm$ 0.0363
-UNI [15] 0.6684 $\pm$ 0.0432 0.5865 $\pm$ 0.0254
-HEST-FT [30] 0.6531 $\pm$ 0.0453 0.5728 $\pm$ 0.0321
-BLEEP [73] 0.6716 $\pm$ 0.0394 0.5778 $\pm$ 0.0383
-RankByGene 0.6814 $\pm$ 0.0512 0.5945 $\pm$ 0.0293
-Table 1: Results for the tasks of gene expression prediction and survival analysis using different patch encoders.
-We use Breast-ST1, Breast-ST2, and Lung-ST datasets for gene expression prediction.
-The slide-level prediction is performed using the ABMIL [28] framework.
-We use UNI [15] as the backbone for the image encoder, followed by a 3-layer MLP as the projection head.
-A 3-layer MLP is also employed for the gene encoder, consistent with the previous method [73].
-Both image and gene feature dimensions are set to 1024.
-For the image encoder, the learning rate is set to 0.0001, while for the gene encoder, it is set to 0.001.
-The other training settings and test setting are detailed in the supplementary materials.
-4.4 Performance Comparison
-We compare our method with image-only pretrained baselines, including ResNet-50 [23] (pretrained on ImageNet), CTransPath [70] and UNI [15] (both pretrained on pathology images), as well as gene-guided image feature learning baselines HEST-FT [30] and BLEEP [73].
-Gene Prediction.
-As shown in Figure 4(a), RankByGene outperforms all baseline methods across the Breast-ST1, Breast-ST2, and Lung-ST datasets.
-Specifically, RankByGene achieves $11\%$ to $42\%$ improvement in PCC over the best baseline performance in each dataset.
-These results indicate that our method achieves stronger alignment between the image and gene expression data, with the obtained gene-informed image features effectively capturing the underlying global relationships between spots.
-Survival Analysis.
-Using the models trained on Breast ST and Lung ST datasets, we evaluate their performance on survival prediction using TCGA-BRCA and TCGA-LUAD WSIs, respectively.
-Figure 4(b) presents the performance of different methods on these two datasets.
-We use the image encoder from each method to extract features on patches, then apply the ABMIL [28] framework to aggregate all patch-level features for slide-level prediction.
-The table shows that RankByGene achieves the highest C-Index of 0.6814 on the TCGA-BRCA dataset and 0.5945 on the TCGA-LUAD dataset compared to all baseline methods.
-The strong performance on these two datasets suggests that the improved image-gene alignment enables the model to capture survival-related signals more effectively, highlighting the potential of our method to enhance performance on challenging downstream tasks.
-4.5 Ablation Study
+ResNet-50 [25] 0.586 $\pm$ 0.052 0.563 $\pm$ 0.044
+CTransPath [74] 0.641 $\pm$ 0.047 0.574 $\pm$ 0.036
+UNI [15] 0.668 $\pm$ 0.043 0.587 $\pm$ 0.025
+HEST-FT [32] 0.653 $\pm$ 0.045 0.573 $\pm$ 0.032
+BLEEP [77] 0.672 $\pm$ 0.039 0.578 $\pm$ 0.038
+RankByGene 0.681 $\pm$ 0.051 0.595 $\pm$ 0.029
+Table 2: Classification and survival analysis results on BCNB and TCGA Datasets. Slide-level prediction is performed with ABMIL [30].
 $\lambda_{1}$ Breast-ST1
 MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
-0 0.3442 0.2516 0.1542
-2 0.3431 0.2499 0.1776
-5 0.3412 0.2487 0.1803
-10 0.3427 0.2505 0.1762
+0 0.3603 0.2751 0.1568
+2 0.3563 0.2735 0.1842
+5 0.3542 0.2723 0.1874
+10 0.3579 0.2744 0.1819
 $\lambda_{2}$ Breast-ST1
 MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
-0 0.3455 0.2535 0.1362
-0.5 0.3438 0.2512 0.1509
-1 0.3412 0.2487 0.1803
-2 0.3436 0.2505 0.1544
+0 0.3611 0.2759 0.1493
+0.5 0.3584 0.2747 0.1602
+1 0.3542 0.2723 0.1874
+2 0.3592 0.2749 0.1587
 $\mathcal{L}_{\text{rank}}$ $\mathcal{L}_{\text{distil}}$ Breast-ST1
 MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
-0.3469 0.2552 0.1269
-0.3455 0.2535 0.1362
-0.3442 0.2516 0.1542
-0.3412 0.2487 0.1803
+0.3624 0.2773 0.1419
+0.3611 0.2759 0.1493
+0.3603 0.2751 0.1568
+0.3542 0.2723 0.1874
 Encoder Breast-ST1
 MAE $\downarrow$ MSE $\downarrow$ PCC $\uparrow$
-ResNet50 0.3633 0.2716 0.0867
-ResNet50 + Ours 0.3573 0.2682 0.1035
-UNI 0.3559 0.2648 0.1063
-UNI + Ours 0.3412 0.2487 0.1803
-Table 2: RankByGene ablation experiments. We conduct all ablation experiments on the gene expression prediction task and the Breast-ST1 dataset.
-Weight of Cross-Modal Ranking Consistency Loss.
-We compare the effect of different values for the weight of the ranking consistency loss $\lambda_{1}$ on the Breast-ST dataset.
-The results in Figure 5(a) show that the model is robust to the choice of the hyperparameter.
-The performance across all the tested values $\lambda_{1}>0$ significantly outperforms the baselines, with the best result using $\lambda_{1}=5$.
-Weight of Intra-Modal Distillation Loss.
-Figure 5(b) presents the ablation study results for the image consistency loss weight $\lambda_{2}$ on the Breast-ST1 dataset.
-Our method shows stable performance across MAE, MSE, and PCC at various values of $\lambda_{2}$, with consistent improvements over the baseline methods.
-These results indicate the benefits of incorporating image consistency loss, as it further enhances the robustness of image feature learning during alignment with gene features.
-Ablation Study on Loss Component.
-Figure 5(c) presents the ablation study results on the impact of individual loss components on the Breast-ST1 dataset.
-When both loss components are removed, the model exhibits the lowest PCC and higher MAE and MSE values.
-Including either $\mathcal{L}_{\text{rank}}$ or $\mathcal{L}_{\text{distil}}$ enhances performance, with the best results achieved when both components are applied.
-The results indicate that both the ranking loss and the image consistency loss contribute significantly to the model’s performance.
-Figure 5(d) presents the results of applying our method to different image encoders on the Breast ST1 dataset, specifically ResNet50 (a CNN-based model pretrained on ImageNet) and UNI (a ViT-based model pretrained on pathology images using the DINO framework [11]).
-The results demonstrate that our method consistently enhances the performance of each encoder in the gene prediction task, as evidenced by lower MAE and MSE values and higher PCC scores.
-This finding shows the adaptability of our approach, demonstrating its effectiveness across various encoder architectures.
+ResNet50 0.3812 0.2914 0.0891
+ResNet50 + Ours 0.3709 0.2833 0.1123
+UNI 0.3692 0.2821 0.1163
+UNI + Ours 0.3542 0.2723 0.1874
+Table 3: Ablation studies. We conduct all ablation experiments on the gene expression prediction task on Breast-ST1 dataset.
+Implementation details.
+In the training stage, We use UNI [15] as the backbone for the image encoder, followed by a 3-layer MLP as the projection head.
+A 3-layer MLP is also employed for the gene encoder, the same as the previous method [77].
+Following [77], we use cosine similarity to calculate the distance between two modalities.
+In the testing stage, we perform 5-fold standalone cross-validation for the gene prediction task.
+The training set is split into five folds, and the best model from each fold is evaluated on the test set.
+For classification and survival prediction tasks, we conduct standard 5-fold cross-validation and report both the mean and standard deviation.
+During inference, we use the teacher model as it performs better.
+Other training settings, test settings, and resource requirements are provided in the supplementary materials.
+Cancer-specific gene selection.
+Since ST gene expression is high dimensional and extremely sparse, existing methods typically select highly expressed genes for learning and prediction, e.g., top 250 [17], top 100 [12], or top 50 [77, 32, 17] genes.
+While these highly expressed genes ensures better quality gene features, we are concerned that they may not be relevant to the biology of interest, and thus may be suboptimal for our downstream tasks.
+Consulting pathologists, we select prognosis-related genes from the Human Protein Atlas [4] as our cancer-specific gene lists.
+This resulted in 447 genes for breast cancer and 1916 genes for lung cancer.
+These gene lists were used for training and evaluation for downstream tasks including gene mutation prediction, receptor status classification, and survival analysis.
+Despite the concern than some of the genes may be close to zero, we observe these cancer-specific genes provide better performance in downstream tasks.
+We stress that this should be adapted as the community moves forward with ST data analysis.
+Please refer to supplementary for further details and discussion.
+When evaluating on gene prediction task, it is important to avoid sparse genes.
+We evaluated on both top 250 highly expressed genes overall (as in [17]) and top 250 highly expressed genes within our cancer-specific gene lists.
+4.1 Quantitative Results
+Figure 5: Comparison of gene-image distances from 100 randomly sampled spot pairs.
+Each point represents the gene and image distance between two spots.
+A higher $R^{2}$ [52] indicates a stronger linear correlation, suggesting better alignment between gene and image features.
+Figure 6: Visualization of FASN gene expression predictions from different methods, with all values normalized to the range of 0 to 1.
+Figure 7: Rank accuracy for different methods during training.
+Gene Expression Prediction.
+As shown in Table 1, RankByGene outperforms all baseline methods across the Breast-ST1, Breast-ST2, and Lung-ST datasets.
+Specifically, RankByGene achieves a PCC improvement of $9\%$ to $34\%$ for the top 250 highly expressed genes and $7\%$ to $35\%$ for the top 250 cancer-specific genes compared to the best baseline performance in each dataset.
+We also provide the results evaluated on the full cancer-specific gene list in the supplementary.
+These results indicate that our method achieves a stronger alignment between image and gene expression data, with the learned gene-informed image features effectively capturing the underlying global relationships between spots.
+Gene-related Classification.
+Using models trained on Breast ST and Lung ST datasets, we evaluate their performance on BCNB and TCGA-LUAD mutation datasets, respectively, as shown in Figure 4(a).
+Each method’s image encoder extracts patch-level features, which are then aggregated using the ABMIL [30] for slide-level prediction.
+RankByGene achieves strong AUC performance on most mutation statuses, demonstrating that enhanced image-gene alignment enables the model to capture gene-related information more effectively.
+This highlights the advantage of integrating gene information for downstream tasks.
+Survival Analysis.
+Following the same setup, we evaluate survival prediction on TCGA-BRCA and TCGA-LUAD WSIs.
+Figure 4(b) shows that RankByGene achieves the highest C-Index (0.6814 for TCGA-BRCA and 0.5945 for TCGA-LUAD), further demonstrating its ability to capture survival-related signals and potential for complex downstream tasks.
+4.2 Ablation Study
+In our ablation studies, we systematically assess the influence of each key component and the selected hyperparameters.
+First, we vary the weights of the cross-modal ranking consistency loss and the intra-modal distillation loss to analyze their individual effects on performance.
+We also perform a comprehensive ablation study on the inclusion of each loss component, illustrating the contribution of each module to the overall alignment quality.
+Furthermore, we compare different image encoders, including CNN-based and Transformer-based to verify the generalizability of our method.
+The results are shown in Table 3.
+A detailed examination is provided in the supplementary.
 4.3 Qualitative Result
-Figure 5: Visualization of PTGES3 gene expression predictions from different methods, with all values normalized to the range of 0 to 1.
-Compared to the baseline, our predictions show the closest match to the ground truth.
 Rank Accuracy During Training.
@@
 In Figure 7, we compare the change of the rank accuracy during training between different methods.
-The figure shows that our method achieves a larger improvement in accuracy during training compared to the baseline methods, suggesting that the rank relationships in the gene features are effectively captured by image features during the training process.
-Figure 6: Rank accuracy for different methods during training.
+Additionally, in Figure 5, we randomly select gene-image pairs to examine whether the distance relationships in gene expression values are reflected in image feature distances.
+The two figures together demonstrate that our method achieves a larger improvement in accuracy during training compared to the baseline methods, suggesting that the rank relationships in the gene features are effectively captured by image features.
 Visualization of Cancer Marker Gene.
-To evaluate the accuracy of our model’s predictions for critical genes, we selected PTGES3 as the marker gene for visualization on the Breast-ST1 dataset due to its association with poor prognosis in breast invasive carcinoma across two independent cohorts [4].
-We visualized the gene prediction values for the PTGES3 from different methods across each spot, including ground truth values and the corresponding WSI region, as shown in Figure 5.
-Compared to the baseline method, our approach provides more accurate predictions for the PTGES3, demonstrating its effectiveness in capturing important gene expression patterns.
-We also provide visualizations of other marker genes for lung cancer in the supplementary materials.
+To evaluate the accuracy of our model’s predictions for critical genes, we selected FASN [24] as the marker gene for visualization on the Breast-ST1 dataset due to its association with poor prognosis in breast invasive carcinoma across two independent cohorts [4].
+We visualized the gene prediction values for the FASN from different methods across each spot, including ground truth values and the corresponding WSI region, as shown in Figure 6.
+Compared to the baseline method, our approach provides more accurate predictions for the FASN, demonstrating its effectiveness in capturing important gene expression patterns.
+We also provide visualizations of other marker genes for breast and lung cancer in the supplementary materials.
 5 Conclusion
-Our proposed framework establishes a robust and scalable approach for efficiently aligning spatial transcriptomics data with histopathology images, overcoming key challenges in cross-modal alignment.
-We propose a ranking-based mechanism to capture similarity relationships between gene and image features, facilitating alignment that is resilient to distortions and effective across both local and global scales.
-This approach preserves essential spatial and molecular patterns, which is critical for accurately capturing the biological complexity of tissue structures.
-Additionally, our use of a self-supervised knowledge distillation method further enhances the alignment process, enabling the model to adaptively handle the high dimensionality, sparsity, and noise inherent in gene expression data.
-Through comprehensive experiments, we demonstrate our framework not only improves alignment quality but also enhances predictive capabilities in gene expression prediction and survival analysis.
-Ultimately, this work provides a foundational tool for multi-modal representation learning in digital pathology.
+We present a robust and scalable approach for aligning spatial transcriptomics data with histopathology images, addressing major obstacles in cross-modal alignment.
+This method employs a ranking-based mechanism to capture similarity relationships between gene and image features, ensuring resilience to distortions and effectiveness across local and global scales while preserving crucial spatial and molecular patterns.
+A self-supervised knowledge distillation procedure further refines alignment by handling high dimensionality, sparsity, and noise in gene expression data.
+Through extensive experiments, our framework not only improves alignment quality, but also enhances predictive performance in gene expression prediction, slide-level classification, and survival analysis.
+This work thus provides a foundational tool for advancing multi-modal representation learning in digital pathology.